# Effective Community Detection Over Streaming Bipartite Networks (Technical Report)

### v1

@@ 13 $\Delta R=\emptyset$;
16
17remove duplicate communities $g$ from $\Delta R$ (overlapping with $R_{t}$) and refine candidate communities $g$ in $\Delta R$;
@@ 18 $R_{t}=R_{t}\cup\Delta R$;
Edge Expiration Processing: To deal with the expiration of an item $p_{t-s}$ ($=(e_{u_{i},v_{a}},t-s)$) in the sliding window $W_{t-1}$, the weight $w_{u_{i},v_{a}}$ of the edge $e_{u_{i},v_{a}}$ decreases (which might turn to the edge deletion). In this case, we will identify those community answers $g$ in the answer set $R_{t}$ that contain edge $e_{u_{i},v_{a}}$ (lines 2-3), and update the community $g$ with $g^{\prime}$ (upon weight/edge changes; line 4). If such a community $g^{\prime}$ does not exist (violating the CD-SBN predicates), we simply remove community answer $g$ from $R_{t}$ (lines 5-6).
Edge Insertion Processing: Next, for the insertion of a new item $p_{t}$ ($=(e_{u_{i}^{\prime},v_{a}^{\prime}},t)$), we aim to update $R_{t}$ with new CD-SBN community answers (containing the new/updated edge $e_{u_{i}^{\prime},v_{a}^{\prime}}$) in a set $\Delta R$ (lines 7-11). In particular, for each user vertex $u_{j}$ within the $2r$-hop away from $u_{i}^{\prime}$, we obtain a CD-SBN candidate community $g$ and add it to $\Delta R$ (if $g$ exists) (lines 8-9). Then, we will remove the redundancy of community candidates in $\Delta R$ (i.e., already in $R_{t}$) and refine the remaining candidate communities in $\Delta R$ (line 10). Finally, we add CD-SBN answers in $\Delta R$ to $R_{t}$, and return final CD-SBN answers in $R_{t}$ (lines 11-12).
Complexity Analysis: Let $deg_{avg}$ be the average degree of the item vertices. For edge insertion and expiration, the time cost of user relationship score re-computation is $O(deg_{avg})$ and of edge support update is $O(deg_{avg}^{2})$. Since obtaining the maximal $(k,r,\sigma)$-bitruss with computed edge supports and user relationship scores (Sarıyüce and Pinar, 2018) requires $O(deg_{avg}^{2})$, it takes $O(|R_{t-1}|deg_{avg}^{2})$ to maintain the result set at $t-1$ and $O(deg_{avg}^{3})$ to detect potential communities. Therefore, the total time complexity of Algorithm 3 at each timestamp $t$ is given by $O\left(|R_{t-1}|deg_{avg}^{2}+deg_{avg}^{3}\right)$.
7. Experimental Evaluation
7.1. Experimental Settings
We conduct a set of experiments to examine the performance of our snapshot and continuous CD-SBN query processing algorithms over different real-world bipartite graphs. Our source code is available on GitHub [^1] https://github.com/L1ANLab/CD-SBN.
@@ 7.1. Experimental Settings
dBkU, dBkL, dBkP Synthetic 25,000 25,000 152,175 N/A
dPkU, dPkL, dPkP Synthetic 25,000 25,000 152,922 N/A
Bipartite Graph Data Sets: We evaluate our proposed CD-SBN algorithms on nine real-world and six synthetic graphs. In particular, as depicted in Table 2, we provide statistics of real-world bipartite graphs from the KONECT [^2] http://konect.cc/ project (i.e., Koblenz Network Collection), where $|\Sigma|$ refers to the domain size of keywords in item vertices of the original graphs (“N/A” for $|\Sigma|$ means that graph data have no keywords).
@@
Competitors: Since no prior works studied the CD-SBN problem under the same semantics as $(k,r,\sigma)$-bitruss communities with specific keywords, we compare our CD-SBN algorithm with a baseline method, named Bitrussness-Based Decomposition (BBD). Specifically, BBD first calculates the trussness of each edge, which is the maximum $k$ of all the $k$-bitruss containing the edge and builds a synopsis based on the trussness. Then, upon edge updates at each timestamp, BBD filters out user vertices connected to edges with trussness less than $k$ and computes the $(k,r,\sigma)$-bitruss community centered at each remaining vertex (containing query keywords).
@@
Measures: We evaluate our CD-SBN performance, in terms of the wall clock time, which is the time cost to traverse the index and retrieve communities for snapshot CD-SBN, or that to maintain the CD-SBN results at each timestamp for continuous CD-SBN problem. The wall clock time is an average time cost over 10 runs with a set of 10 different CD-SBN predicates (e.g., query keywords).
(a) real-world graphs
(b) synthetic graphs
Figure 3. CD-SBN performance on data sets.
Parameter Settings: Table 3 depicts the parameter settings of our subsequent experiments, where default values are in bold. Each time we vary one parameter while setting other parameters to their default values. Moreover, since our continuous CD-SBN processing algorithm has no data dependencies during the search of communities, we use multithreaded optimization to speed up the computation, where the number of threads is set to $20$ by default. All the experiments are executed on a machine with an AMD Ryzen Threadripper 3990X 64-Core processor, Ubuntu 22.04 OS, 192GB memory, and a 600GB disk. All algorithms were implemented in C++ (C++17 standard) and compiled with the g++ 11.4.0 compiler.
RQ1 (Efficiency): Can our proposed algorithms efficiently process continuous CD-SBN problem?
@@
Figure 4. Robustness evaluation of CD-SBN.
@@ 7.2. CD-SBN Performance Evaluation
We also test the robustness of our CD-SBN approach with different parameters (e.g., $s$, $k$, $r$, $\sigma$, and $|Q|$) on synthetic graphs with different structural parameters (e.g., $|\Sigma|$, $|v_{i}.K|$, $[min\_w,max\_w]$, $|L(G)|$, and $|U(G)|$).
@@
Effect of Sliding Window Size $s$: Figure 4(a) presents the performance of our continuous CD-SBN approach, by varying the size $s$ of sliding window $W_{t}$ from $50$ to $200$, where other parameters are by default. From the experimental results, with the increasing $s$, more edges are involved or have higher edge weights, which result in more potential candidate communities. Therefore, larger window size $s$ incurs higher time cost. Nonetheless, the time cost of our CD-SBN approach remains low (i.e.,0.047 $\sim$ 0.726 $sec$) for different $s$ values.
@@
Effect of Bitruss Support Threshold $k$: Figure 4(b) illustrates the continuous CD-SBN performance, where the bitruss support parameter $k$ varies from $3$ to $5$, and other parameters are set to default values. Since a larger support threshold $k$ leads to higher pruning power, our CD-SBN approach can obtain fewer candidate communities, which in turn incurs lower time costs. The query cost remains low for different $k$ values (i.e., 0.053 $\sim$ 0.277 $sec$).
@@
Effect of Radius $r$: Figure 4(c) shows the experimental results of our CD-SBN approach, by varying radius, $r$, the query community from $1$ to $3$, where default values are used for other parameters. When radius $r$ becomes larger, the scale of $2r$-hop subgraphs will increase, which results in higher time costs to obtain maximal $(k,r,\sigma)$-bitruss. Nevertheless, for different $r$ values, the wall clock time remains small (i.e., around 0.033 $\sim$ 0.375 $sec$).
@@
Effect of Relationship Score Threshold $\sigma$: Figure 4(d) varies the relationship score threshold, $\sigma$, from $1$ to $5$, where other parameters are set by default values. Since higher $\sigma$ leads to higher pruning power and cohesiveness of candidate communities, the number of candidate communities decreases for larger $\sigma$. Thus, the CD-SBN time cost decreases for larger $\sigma$ values. Overall, for different $\sigma$ values, the time cost remains low (i.e., 0.051$\sim$1.989 $sec$).
@@
Effect of Query Keyword Set Size $|Q|$: Figure 4(e) evaluates the CD-SBN performance, with different sizes, $|Q|$, of the query keyword set $Q$ from $2$ to $10$, where we use default values for other parameters. Intuitively, a larger query keyword set $Q$ allows more item vertices to be included (with lower pruning power by the keyword pruning), and in turn more candidate communities will be retrieved for filtering/refinement. Thus, from the figure, when $|Q|$ becomes larger, the query cost also increases. Nonetheless, the wall clock time remains low (i.e., 0.007 $\sim$ 0.893 $sec$) for different $|Q|$ values.
@@
Effect of Keyword Domain Size $|\Sigma|$: Figure 4(f) shows the wall clock time of our CD-SBN approach, where the keyword domain size, $|\Sigma|$, varies from $100$ to $1,000$ and other parameters are set to their default values. In the figure, we can see that, when $|\Sigma|$ becomes larger, our CD-SBN approach requires less time to retrieve CD-SBN communities, which indicates the effectiveness of our keyword pruning for large keyword domain size. Overall, with different $|\Sigma|$ values, the time cost remains low (i.e., 1.974 $\sim$ 0.009 $sec$).
@@
Effect of # of Keywords Per Item Vertex $|v_{i}.K|$: Figure 4(g) illustrates the CD-SBN performance for different numbers, $|v_{i}.K|$, of keywords per item vertex from $1$ to $5$, where we set other parameters to default values. Intuitively, more keywords in item vertices $v_{i}$ will increase the chance to match with query keyword set $Q$, which results in more or larger candidate communities to filter and refine. Therefore, with the increase of $|v_{i}.K|$, the time cost of our CD-SBN approach also increases, but, nevertheless, remains low (i.e., 0.014$\sim$0.821 $sec$).
Effect of Edge Weight Range $\left[min\_w,max\_w\right]$: Figure 4(h) examines the CD-SBN performance with different edge weight ranges, $\left[min\_w,max\_w\right]$ (i.e., $[1,2]$, $[1,3]$, and $[1,4]$), where default values are used for other parameters. When the range of edge weights expands, more/larger candidate communities will be retrieved and refined. Thus, the query cost increases as the edge weight interval $\left[min\_w,max\_w\right]$ becomes wider. Nonetheless, the wall clock time remains low (i.e., 0.0543$\sim$2.761 $sec$).
Effect of Lower Layer Size $|L(G)|$: Figure 4(i) tests the scalability of our proposed CD-SBN approach by varying the lower layer size, $|L(G)|$, from $10K$ to $100K$, where default values are used for other parameters. From the figure, we can see that the query cost decreases as $|L(G)|$ increases. This is because, with a constant average degree of each user vertex, larger $|L(G)|$ leads to a more sparse bipartite network, which incurs fewer candidate communities and lower query costs. The time cost remains low (i.e., 0.031$\sim$0.407 $sec$).
@@
Effect of Upper Layer Size $|U(G)|$: Figure 4(j) evaluates the scalability of our CD-SBN approach with different upper layer sizes, $|U(G)|$, from $10K$ to $100K$, where other parameters are set by default. In the figure, with the increase of $|U(G)|$, the CD-SBN time cost increases due to more retrieved candidate communities. Nonetheless, the time cost remains low (i.e., 0.001 $\sim$ 1.815 $sec$), which confirms the good scalability of our CD-SBN approach for large $|U(G)|$.
@@
Ablation Study on the Snapshot CD-SBN Pruning Strategies (RQ2): To test the effectiveness of our proposed pruning methods, we conduct an ablation study of our snapshot CD-SBN pruning strategies over real-world/synthetic graphs, where all parameters are set to default values. We execute the snapshot CD-SBN processing algorithm by adding one pruning strategy at a time, which forms 3 experimental groups: (1) support pruning only, (2) support + score pruning, and (3) support + score + keyword pruning. Figure 5 reports the number of pruned candidate communities by using different pruning method groups. In this figure, we can see that for all data sets, the number of pruned communities increases by about an order of magnitude, as we apply more pruning strategies, which indicates the effectiveness of our proposed pruning strategies.
Figure 5. Ablation study of the pruning power for the snapshot CD-SBN problem.
Figure 6. Case study of $(k,r,\sigma)$-bitruss vs. biclique (Lin, 2018).
@@
CD-SBN Community Case Study (RQ3): To evaluate the utility of our CD-SBN results, we carry out a case study to compare our proposed $(k,r,\sigma)$-bitruss pattern with the biclique (Lin, 2018) (which is a variant of bipartite community structure) over BibSonomy (BS) graph data. The entire graph shown in Figure 6 is a $(4,2,3)$-bitruss containing 3 users (in blue) and 9 items (in blue or red), whereas the 3 users, 4 items (in blue) form a $3\times 4$ biclique. We can see that the $3\times 4$ biclique shares the same 3 users as that in $(4,2,3)$-bitruss, but the $3\times 4$ biclique ignores some items that have edges connected to some users with large weights (e.g., 6). In contrast, our $(k,r,\sigma)$-bitruss captures these items, and can recommend them to users who are not connected to them (e.g., recommending items $v\_449272$, $v\_449273$, $v\_449277$, $v\_449275$ and $449276$ to user $u\_4954$). This confirms the utility of our proposed CD-SBN community semantics, which can recommend items to more potentially interested users.

### v2

@@ 12 $\Delta R=\emptyset$;
15
16remove duplicate communities $g$ from $\Delta R$ (overlapping with $R_{t}$) and refine candidate communities $g$ in $\Delta R$;
@@ 17 $R_{t}=R_{t}\cup\Delta R$;
6.3. Continuous CD-SBN Processing Algorithm
Algorithm 3 details our continuous CD-SBN query processing approach. First, we initialize the CD-SBN result set $R_{t}$ at timestamp $t$ with that $R_{t-1}$ at timestamp (t1)(t-1) (line 1). Next, we check the constraints of the existing communities to refine the previous result set (lines 2-6). Then, we obtain the potential communities from the $2r$-hop of user vertices near the insertion edge and remove duplicates (lines 7-10). Finally, we merge the candidate communities into the result set and return it at timestamp $t$ (lines 11-12).
Edge Expiration Processing: To deal with the expiration of an item $p_{t-s}$ ($(e_{u_{i},v_{a}},t-s)$) in the sliding window $W_{t-1}$, the weight $w_{u_{i},v_{a}}$ of the edge $e_{u_{i},v_{a}}$ decreases (which might turn to the edge deletion). In this case, we will identify those community answers $g$ in the answer set $R_{t}$ that contain edge $e_{u_{i},v_{a}}$ (lines 2-3), and update the community $g$ with $g^{\prime}$ (upon weight/edge changes; line 4). If such a community $g^{\prime}$ does not exist (violating the CD-SBN predicates), we simply remove community answer $g$ from $R_{t}$ (lines 5-6).
Edge Insertion Processing: Next, for the insertion of a new item $p_{t}$ ($(e_{u_{i}^{\prime},v_{a}^{\prime}},t)$), we aim to update $R_{t}$ with new CD-SBN community answers (containing the new/updated edge $e_{u_{i}^{\prime},v_{a}^{\prime}}$) in a set $\Delta R$ (lines 7-11). In particular, for each user vertex $u_{j}$ within the $2r$-hop away from $u_{i}^{\prime}$, we obtain a CD-SBN candidate community $g$ via the pruning strategies mentioned in Section 4 and add it to $\Delta R$ (if $g$ exists) (lines 8-9). Then, we will remove redundancy among community candidates in $\Delta R$ (i.e., those already in $R_{t}$) and refine the remaining candidate communities in $\Delta R$ (line 10). Finally, we add CD-SBN answers in $\Delta R$ to $R_{t}$, and return final CD-SBN answers in $R_{t}$ (lines 11-12).
Complexity Analysis: Let $deg_{avg}$ be the average degree of the item vertices. For edge insertion and expiration, the time cost of user relationship score re-computation is $O(deg_{avg})$ and of edge support update is $O(deg_{avg}^{2})$. Since obtaining the maximal $(k,r,\sigma)$-bitruss with computed edge supports and user relationship scores (Sarıyüce and Pinar, 2018) requires $O(deg_{avg}^{2})$, it takes $O(|R_{t-1}|deg_{avg}^{2})$ to maintain the result set at t1t-1 and $O(deg_{avg}^{3})$ to detect potential communities. Therefore, the total time complexity of Algorithm 3 at each timestamp $t$ is given by $O\left(|R_{t-1}|deg_{avg}^{2}+deg_{avg}^{3}\right)$.
@@ 6.3. Continuous CD-SBN Processing Algorithm
[…]
7. Experimental Evaluation
7.1. Experimental Settings
We conduct experiments to evaluate the performance of our snapshot and continuous CD-SBN query processing algorithms on various bipartite graphs. Our source code is available on GitHub [^1] https://github.com/L1ANLab/CD-SBN.
Bipartite Graph Data Sets: We evaluate our proposed CD-SBN algorithms on nine real-world and six synthetic graphs. As depicted in Table 2, we provide statistics of real-world bipartite graphs from the KONECT [^2] http://konect.cc/ project (i.e., Koblenz Network Collection), where $|\Sigma|$ refers to the domain size of keywords in item vertices of the original graphs (“N/A” for $|\Sigma|$ means that graph data have no keywords).
@@ 7.1. Experimental Settings
Competitors: Since no prior work has studied the CD-SBN problem under the same semantics as $(k,r,\sigma)$-bitruss communities with specific keywords, we compare our CD-SBN algorithm with a baseline method, Bitrussness-Based Decomposition (BBD). Specifically, BBD first calculates the trussness of each edge, which is the maximum $k$ of all the $k$-bitruss containing the edge, and builds a synopsis based on the trussness. Then, upon edge updates at each timestamp, BBD filters out user vertices connected to edges with trussness less than $k$ and computes the $(k,r,\sigma)$-bitruss community centered at each remaining vertex (containing query keywords).
Measures: We evaluate our CD-SBN performance in terms of the wall clock time, which is the time cost to traverse the index and retrieve communities for snapshot CD-SBN, or that to maintain the CD-SBN results at each timestamp for the continuous CD-SBN problem. The wall clock time is the average time across 10 runs with a set of 10 different CD-SBN predicates (e.g., query keywords).
@@
Figure 3. The robustness evaluation of the CD-SBN approach.:
Parameter Settings: Table 3 depicts the parameter settings of our subsequent experiments, where default values are in bold. Each time we vary one parameter while keeping the others at their default values. Moreover, we use multithreaded optimization to speed up the computation of candidates, where the number of threads is set to $100$ by default. All experiments are conducted on a machine equipped with an AMD Ryzen Threadripper 3990X 64-Core processor, running Ubuntu 22.04 OS, 192GB of memory, and a 600GB disk. All algorithms were implemented in C++ (C++17 standard) and compiled with g++ 11.4.0.
RQ1 (Efficiency): Can our proposed algorithms efficiently process the continuous CD-SBN problem?
@@ 7.2. CD-SBN Performance Evaluation
We also test the robustness of our CD-SBN approach with different parameters (e.g., $s$, $k$, $r$, $\sigma$, and $|Q|$) on synthetic graphs with different structural parameters (e.g., $|\Sigma|$, $|v_{i}.K|$, [min_w,max_w][min\_w,max\_w], $|L(G)|$, and $|U(G)|$).
@@
Effect of Sliding Window Size $s$: Figure 3(a) presents the performance of our continuous CD-SBN approach, by varying the size $s$ of the sliding window $W_{t}$ from $50$ to $200$, where other parameters are by default. From the experimental results, as $s$ increases, more edges are involved or have higher edge weights, resulting in more potential candidate communities. Therefore, a larger window size $s$ incurs a higher time cost, but the time cost of our CD-SBN approach remains low (i.e.,0.047 $\sim$ 0.726 secsec).
@@
Effect of Bitruss Support Threshold $k$: Figure 3(b) illustrates the continuous CD-SBN performance, where the bitruss support parameter $k$ varies from $3$ to $5$, and other parameters are set to default values. Since a larger support threshold $k$ yields greater pruning power, our CD-SBN approach can identify fewer candidate communities, thereby incurring lower time costs. The query cost remains low for different $k$ values (i.e., 0.055 $\sim$ 0.249 secsec).
(a) real-world graphs
(b) synthetic graphs
Figure 4. The CD-SBN performance on real/synthetic graphs.:
@@
Effect of Radius $r$: Figure 3(c) shows the experimental results of our CD-SBN approach, by varying radius, $r$, the query community from $1$ to $3$, where default values are used for other parameters. When the radius $r$ becomes larger, the scale of $2r$-hop subgraphs will increase, which results in higher time costs to obtain maximal $(k,r,\sigma)$-bitruss. Nevertheless, for different $r$ values, the wall clock time remains small (i.e., around 0.032 $\sim$ 0.256 secsec).
@@
Effect of Relationship Score Threshold $\sigma$: Figure 3(d) varies the relationship score threshold, $\sigma$, from $1$ to $5$, where other parameters are set by default values. Since higher $\sigma$ leads to higher pruning power and cohesiveness of candidate communities, the number of candidate communities decreases for larger $\sigma$. Thus, the CD-SBN time cost decreases for larger $\sigma$ values. Overall, for different $\sigma$ values, the time cost remains low (i.e., 0.059 $\sim$ 0.697 secsec).
@@
Effect of Query Keyword Set Size $|Q|$: Figure 3(e) evaluates the CD-SBN performance, with different sizes, $|Q|$, of the query keyword set $Q$ from $2$ to $10$, where we use default values for other parameters. Intuitively, a larger set of query keywords $Q$ allows more item vertices to be included, resulting in lower pruning power by the keyword pruning. Consequently, this leads to the retrieval of more candidate communities for further filtering and refinement. Thus, from the figure, as $|Q|$ increases, the query cost also increases. Nonetheless, the wall clock time remains low (i.e., 0.007 $\sim$ 0.893 secsec) for different $|Q|$ values.
@@
Effect of Keyword Domain Size $|\Sigma|$: Figure 3(f) shows the wall clock time of our CD-SBN approach, where the keyword domain size, $|\Sigma|$, varies from $100$ to $1,000$ and other parameters are set to their default values. In the figure, we find that as $|\Sigma|$ increases, our CD-SBN approach requires less time to retrieve CD-SBN communities, indicating the effectiveness of our keyword pruning for large keyword domains. Overall, with different $|\Sigma|$ values, the time cost remains low (i.e., 1.761 $\sim$ 0.021 secsec).
@@
Effect of # of Keywords Per Item Vertex $|v_{i}.K|$: Figure 3(g) illustrates the CD-SBN performance for different numbers, $|v_{i}.K|$, of keywords per item vertex from $1$ to $5$, where we set other parameters to default values. Intuitively, more keywords in item vertices $v_{i}$ will increase the chance to match with query keyword set $Q$, which results in more/larger candidate communities to filter and refine. Therefore, with the increase of $|v_{i}.K|$, the time cost of our CD-SBN approach also increases, but, nevertheless, remains low (i.e., 0.028 $\sim$0.773 secsec).
[…]
Figure 5. The ablation study of the pruning power for the snapshot CD-SBN problem.:
@@
Effect of Upper Layer Size $|U(G)|$: Figure 3(j) evaluates the scalability of our CD-SBN approach with different upper layer sizes, $|U(G)|$, from $10K$ to $100K$, where other parameters are set by default. In the figure, as $|U(G)|$ increases, the CD-SBN time cost increases due to more retrieved candidate communities. Nonetheless, the time cost remains low (i.e., 0.001 $\sim$ 1.815 secsec), which confirms the good scalability of our CD-SBN approach for large $|U(G)|$.
@@
Ablation Study on the Snapshot CD-SBN Pruning Strategies (RQ2): To test the effectiveness of our proposed pruning methods, we conduct an ablation study of our snapshot CD-SBN pruning strategies on real-world/synthetic graphs, with all parameters set to default values. We execute the snapshot CD-SBN processing algorithm by adding one pruning strategy at a time, which forms 3 experimental groups: (1) support pruning only, (2) support + score pruning, and (3) support + score + keyword pruning. Figure 5 reports the number of pruned candidate communities by using different pruning method groups. In this figure, we observe that the number of pruned communities increases by approximately an order of magnitude as we apply more pruning strategies, indicating the effectiveness of our proposed strategies.
@@
CD-SBN Community Case Study (RQ3): To evaluate the utility of our CD-SBN results, we carry out a case study to compare our proposed $(k,r,\sigma)$-bitruss pattern with the biclique (Lin, 2018) (which is a variant of bipartite community structure) over BibSonomy (BS) graph data. The entire graph shown in Figure 6 is a $(4,2,3)$-bitruss containing 3 users (in blue) and 9 items (in blue or red), whereas the 3 users, 4 items (in blue) form a $3\times 4$ biclique. We can see that the $3\times 4$ biclique shares the same 3 users as the $(4,2,3)$-bitruss, but the $3\times 4$ biclique omits some items that are connected to users with large weights (e.g., 6). In contrast, our $(k,r,\sigma)$-bitruss captures these items, and can recommend them to users who are not connected to them (e.g., recommending items v_449272v\_449272, v_449273v\_449273, v_449277v\_449277, v_449275v\_449275, and $449276$ to user u_4954u\_4954). This confirms the utility of our proposed CD-SBN community semantics, which can recommend items to users who are potentially more interested.